OpenStudy (anonymous):

Will fan and medal!!!! Find the product. Then simplify your answer. 9s/-7 x 1/3s

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

\[\frac{9x}{-7}\times \frac{1}{3s}\] like that ??

OpenStudy (anonymous):

Yes

OpenStudy (misty1212):

\[\frac{9s}{-7}\times \frac{1}{3s}\]

OpenStudy (anonymous):

its 9s though

OpenStudy (misty1212):

cancel the threes and also the s

OpenStudy (misty1212):

\[\frac{9s}{-7}\times \frac{1}{3s}=\frac{9\cancel s}{-7}\times \frac{1}{3\cancel s}\]\[=\frac{9}{-7\times 3}=\frac{\cancel{9}^3}{-7\cancel3}\]\[=-\frac{3}{7}\]

OpenStudy (anonymous):

Thank you so much :)